# Density-Matrix Mean-Field Theory

### v1

@@ 4 Discussions
In Sec. 4.1, we undertake a comparative analysis of DMMFT with DMRG and DMFT. Additionally, in Sec.4.2, we extend the DMMFT for systems at finite temperatures, and in Sec. 4.3, we extend the DMMFT for systems with disorders.
@@ 4.1 Comparisons with Alternative Methods
We first compare DMMFT with DMRG. The constructions of the effective Hamiltonian appear similar, where the reduced DM is employed to select Hilbert subspace significant with respect to the target state. The resemblance between the DMMFT and the infinite DMRG is most pronounced when the spatial dimension $d=1$ as illustrated in the AKLT model in Sec. 3.1. The primary difference lies in where the Hilbert space ~$\tilde{\mathscr{H}}=\mathscr{H}_{c}\otimes\tilde{\mathscr{H}}_{\tilde{c}}$ is chosen. In DMMFT, the projector $\Pi_{\mathcal{C}^{\prime}_{c}}$ is iteratively optimized solely over the Hilbert space $\mathscr{H}_{\mathcal{C}^{\prime}_{c}}$ without changing the cluster size of the environment. Conversely, in infinite DMRG, the environment is constructed iteratively over the Hilbert space of the semi-infinite chains. Gauged by the quantum entanglement, the dimension of the Hilbert subspace selected for constructing the effective Hamiltonian is bounded by $\tilde{D}^{E}_{c}$ [Eq.(16)] in both cases. As $\tilde{D}^{E}_{c}$ is bounded by the size of the cluster $c$, DMRG can feasibly select a finite significant Hilbert subspace without being hindered by the infinite size of the environment asymptotically. On the contrary, in DMMFT, we assume that the significant Hilbert subspace resides within $\mathscr{H}_{\mathcal{C}^{\prime}_{c}}$, which is a reasonable approximation, especially for short-range entangled systems. Ignoring practical constraints provided we choose the cluster size being larger than the typical entanglement range. the DMMFT and the DMRG are equivalent in $d=1$.
@@
The advantages of DMMFT become more significant in higher dimensions, $d\geq 2$. In DMRG, a $d$-dimensional system is compactified to a quasi-one-dimensional system, and an artificial one-dimensional chain of clusters winds through the system. Consequently, even in the case of a short-range entangled system, artificial long-range entanglements emerge for sites physically close but distant along the chain. In contrast, DMMFT makes use of the separability [Eq.(17)] enables the study of local problems over extended clusters without introducing artificial long-range entanglements arising from the transverse dimensions. From this perspective, DMMFT may be regarded as a mean-field approximation of DMRG and proves particularly valuable for short-range entangled systems in higher dimensions.
@@
Now, we turn to compare DMMFT with DMFT. In DMFT, designed for fermionic systems, the effective action is constructed over a local cluster, employing the single-particle Green’s function as a dynamical mean-field to capture fluctuations [19, 20]. This local problem, entailing a dynamical mean-field, is subsequently mapped to an effective impurity model and solved using ED or QMC kernels [19]. In contrast to the Anderson impurity model tailored for fermions, establishing a comparable impurity model for spins is not a straightforward task. Furthermore, even if a spin impurity model aligned with the principles of DMFT could be formulated, the challenge persists in efficiently solving it. Constructing a generic spin impurity (if not impossible) can introduce significant frustrations, which leads to the failure of the QMC kernel due to the sign problem. On the other hand, achieving an accurate representation of the environment requires incorporating a large number of spins beyond those in the local cluster, which causes the ED kernel to rapidly encounter the exponential wall. A simplification of DMFT proposed in DMET for fermionic systems involves utilizing the local density matrix, a static observable, instead of the dynamical Green’s function [21]. While a formal generalization of DMET to spin systems is feasible, the challenge lies in identifying a suitable reference wave-function that facilitates straightforward Schmidt decomposition and the construction of the embedding Hamiltonian. On the contrary, in DMMFT, the reduced density matrix is generally defined for many-body states, making DMMFT naturally applicable to fermions, bosons, as well as spins.
@@
Integrating wave function approaches with DMMFT raises the question of determining a global state from reduced density matrices of local clusters. This challenge is recognized as the quantum marginal problem [56]. Finding a general solution to this problem is currently beyond the scope of this paper and remains an open and challenging research question [57].
@@ 4.2 Systems at Finite Temperatures
where $k_{B}$ is the Boltzmann constant, and the square bracket with a subscript $T$ indicates the thermal average. As discussed in Sec. 2.3, when $D_{\mathscr{H}_{\mathcal{C}^{\prime}_{c}}}=1$, DMMFT reduces to the CMFTs at zero temperature. It is straightforward to check that, since $[\rho_{c}^{\text{MF}}]_{T}\sim\exp[-\tilde{H}_{\text{MF}}^{(c)}/(k_{B}T)]$ when $D_{\mathscr{H}_{\mathcal{C}^{\prime}_{c}}}=1$, the extension of DMMFT according Eq.(25) also aligns with the conventional MFTs at finite temperatures.
However, it is worth noting that the thermal averaging in Eq.(25) does not fully capture all thermal fluctuations. Specifically, at low temperatures, it is often the gapless long wave length modes (Goldstone modes) that dominate the thermal fluctuations. However, due to the mean-field approximation inherent in DMMFT, these thermal fluctuations are neglected. Therefore, similar to all other mean-field methods, DMMFT tends to underestimate the impact of thermal fluctuations at finite temperatures.
@@ 4.3 Systems with Disorders
In real systems, the effects of disorders can be more complicated. For instance, there can be vacancies clusters due to lower formation free energy, and the probability measure of the vacancies no longer takes a simple product form. Moreover, disorders can manifest as inhomogeneity of the samples, especially when scale of the probe is small, and the effects of disorders are not self-averaged in the experiments. Despite of the complications of disorders arising in real systems, DMMFT can, in principle, be adapted to include the disorder effects accordingly.
@@ 5 Conclusion
In this study, we introduce a novel mean-field approach, DMMFT, specifically designed to capture quantum fluctuations beyond conventional MFTs. Combining the strengths of DMRG and DMFT, DMMFT stands out as an efficient and unbiased method applicable to fermions, bosons, and spins, even in the presence of frustrations. A noteworthy feature of DMMFT is its ability to discern topological phases, a capability further enhanced by its utilization of the reduced density matrix. Additionally, DMMFT seamlessly extends to systems at finite temperatures and with disorders. In conclusion, DMMFT provides an effective tool for exploring phases exhibiting unconventional quantum orders, particularly beneficial for investigating frustrated spin systems in high spatial dimensions.

### v2

@@ 4 Discussions
In Sec. 4.1, we undertake a comparative analysis of DMMFT with DMRG, DMFT, CVM, and LCE. Additionally, in Sec.4.2, we extend the DMMFT for systems at finite temperatures, and in Sec. 4.3, we extend the DMMFT for systems with disorders.
@@ 4.1 Comparisons with Alternative Methods
4.1.1 Comparison with Density-Matrix Renormalization Group
@@
We first compare DMMFT with DMRG. The constructions of the effective Hamiltonian appear similar, where the reduced DM is employed to select Hilbert subspace significant with respect to the target state. The resemblance between the DMMFT and the infinite DMRG is most pronounced when the spatial dimension $d=1$ as illustrated in the AKLT model in Sec. 3.1. The primary difference lies in where the Hilbert space ~$\tilde{\mathscr{H}}=\mathscr{H}_{c}\otimes\tilde{\mathscr{H}}_{\tilde{c}}$ is chosen. In DMMFT, the projector $\Pi_{\mathcal{C}^{\prime}_{c}}$ is iteratively optimized solely over the Hilbert space $\mathscr{H}_{\mathcal{C}^{\prime}_{c}}$ without changing the cluster size of the environment. Conversely, in infinite DMRG, the environment is constructed iteratively over the Hilbert space of the semi-infinite chains. Gauged by the quantum entanglement, the dimension of the Hilbert subspace selected for constructing the effective Hamiltonian is bounded by $\tilde{D}^{E}_{c}$ [Eq.(16)] in both cases. As $\tilde{D}^{E}_{c}$ is bounded by the size of the cluster $c$, DMRG can feasibly select a finite significant Hilbert subspace without being hindered by the infinite size of the environment asymptotically. On the contrary, in DMMFT, we assume that the significant Hilbert subspace resides within $\mathscr{H}_{\mathcal{C}^{\prime}_{c}}$, which is a reasonable approximation, especially for short-range entangled systems. Ignoring practical constraints provided we choose the cluster size being larger than the typical entanglement range, DMMFT and DMRG are equivalent in $d=1$.
@@
The advantages of DMMFT become more significant in higher dimensions, $d\geq 2$. In DMRG, a $d$-dimensional system is compactified to a quasi-one-dimensional system, and an artificial one-dimensional string of clusters winds through the system as shown in Fig. 3(a). Consequently, even in the case of a short-range entangled system, artificial long-range entanglements emerge for sites physically close but distant along the string. The entanglement between physically neighboring clusters scales exponentially with $N_{\perp}^{d-1}$, where $N_{\perp}$ represents the linear scale of the system along the transverse directions. In contrast, DMMFT makes use of the separability [Eq.(17)] enables the study of local problems over extended clusters without introducing artificial long-range entanglements arising from the transverse dimensions. From this perspective, DMMFT may be regarded as a mean-field approximation of DMRG and proves particularly valuable for short-range entangled systems in higher dimensions.
4.1.2 Comparison with Dynamical Mean-Field Theory
@@
We proceed to compare DMMFT with DMFT. In DMFT, designed for fermionic systems, the effective action is constructed over a local cluster, employing the single-particle Green’s function as a dynamical mean-field to capture the fluctuations [19, 20]. This local problem, entailing a dynamical mean-field, is subsequently mapped to an effective impurity model and solved using ED or QMC kernels [19]. In contrast to the Anderson impurity model tailored for fermions, devising an equivalent impurity model for spins is not straightforward. Furthermore, even if a spin impurity model consistent with the principles of DMFT could be formulated, the challenge of efficiently solving it would remain. Constructing a generic spin impurity (if not impossible) can introduce significant frustrations, which leads to the failure of the QMC kernel due to the sign problem. On the other hand, achieving an accurate representation of the environment requires incorporating a large number of spins beyond those in the local cluster, which causes the ED kernel to rapidly encounter the exponential wall. A simplification of DMFT proposed in DMET for fermionic systems involves utilizing the local density matrix, a static observable, instead of the dynamical Green’s function [21]. While a formal generalization of DMET to spin systems is feasible, the challenge lies in identifying a suitable reference wave-function that facilitates straightforward Schmidt decomposition and the construction of the embedding Hamiltonian. On the contrary, in DMMFT, the reduced DM is generally defined for many-body states, making DMMFT naturally applicable to fermions, bosons, as well as spins.
@@
Moreover, integrating wave function approaches with DMMFT raises the question of determining a global state from reduced density matrices of local clusters. This challenge is recognized as the quantum marginal problem [77]. Finding a general solution to this problem is currently beyond the scope of this paper and remains an open and challenging research question [78].
It deserves particular comments on the lattice symmetries when comparing DMMFT with cluster extensions of DMFT. The problem of the lattice symmetry is a shared problem not only for DMFT but also for all cluster approaches formulated in real space. In cluster extensions of DMFT, this challenge arises because the approximate self-energy might not preserve the underlying lattice symmetry. Inconsistencies may arise if artificial periodization is not appropriately imposed [79]. In contrast, DMMFT works with local static Hamiltonians, which allows for a more straightforward imposition of symmetrization. Let $G={g}$ denote a subgroup of symmetry transformations of the lattice that one would enforce, and denote the representations of $g$ over $\tilde{\mathscr{H}}$ as $\tilde{g}$. Then, the mean-field Hamiltonian in Eq. (11) can be symmetrized as
$\begin{split}\tilde{H}_{\text{MF,Symm}}^{(c)}=\frac{1}{|G|}\sum_{g\in G}\tilde{g}\tilde{H}_{\text{MF}}^{(c)}\tilde{g}^{-1}.\end{split}$ (25)
Given that the reduced density matrix is constructed from the eigenstates of the Hamiltonian, it should naturally respect the corresponding symmetries inherited from the symmetrized Hamiltonian $\tilde{H}_{\text{MF,Symm}}^{(c)}$.
4.1.3 Comparison with Cluster Variation Method
CVM is a variational method minimizing the free energy
$\begin{split}F[\rho]=\text{Tr}(\rho H)+k_{B}T\text{Tr}[\rho\ln(\rho)],\end{split}$ (26)
with respect to the density matrix. In Eq. (26), $\rho$ is the many-body DM for the entire system, $k_{B}$ is the Boltzmann constant, and $T$ is the temperature. For Hamiltonian consisting of terms with finite-range interactions, it is sufficient to know up to $n$-point reduced DM,
$\begin{split}\rho^{(n)}(\{i_{1},i_{2},\dots,i_{n}\})=\text{Tr}_{\mathcal{I}\backslash\{i_{1},i_{2},\dots,i_{n}\}}\rho,\end{split}$ (27)
for evaluating the energetic contribution $\text{Tr}(\rho H)$ to the free energy, where the set $\{i_{1},i_{2},\dots,i_{n}\}$ supports of the terms in the Hamiltonian. For example, in Heisenberg models with exchange interactions and Zeeman couplings, it only requires $\rho^{(n=2)}(i_{1},i_{2})$ where the distance $|i_{1}-i_{2}|$ is bounded by the interaction range of the Hamiltonian. The key idea of the CVM is to use cluster entropies for evaluating the entropic contribution $\text{Tr}[\rho\ln(\rho)]$ to the free energy. The cluster entropies rely only on the reduced DMs supporting the clusters and are expected to decay rapidly when the linear size of the clusters exceeds the correlation length of the system [32]. The cluster approximation involves selecting a truncation in cluster size. A particular choice of the truncation for models with Heisenberg exchange interactions is at second order, i.e., $n=2$, which is also known as the pair approximation [30, 31, 29]. In this approach, the variational equations, along with the consistency conditions of the reduced DM, lead to the effective mean-field Hamiltonians of conventional MFTs [30, 31] and the separability $\rho^{(n=2)}(i,j)=\rho^{(n=1)}(i)\otimes\rho^{(n=1)}(j)$ for $|i-j|$ exceeds the interaction range [29].
The separability in CVM arises as a consequence of the absence of long-range interactions and the truncation at $n=2$. However, quantum fluctuations can manifest as entanglements over longer range, which is inherently determined by the nature of quantum states. In this sense, DMMFT improves over CVM with pair approximation by considering the quantum entanglements over the range of the extended cluster and determining the reduced DM $\rho_{c}$ self-consistently, up to $n=|c|$.
Another question worth mentioning is whether the mean-field equations derived in Sec. 2 can be derived from the perspective of CVM. As demonstrated in Ref.29, conventional mean-field theories [30, 31] can be derived as special cases of CVM with pair approximation and constant-coupling approximation. Different choices of consistency conditions can lead to different conventional mean-field theories [30, 31]. In this article, we have derived the mean-field equations for DMMFT by assuming the separability for distant clusters. However, whether separability can emerge as a natural consequence of alternative approximations beyond pair approximation, as demonstrated for conventional MFTs in CVM in Ref. 29), remains an open question. We do not have an answer to this question and leave it for future research.
4.1.4 Comparison with Linked-Cluster Expansion
LCE is another cluster-based method used for analyzing quantum spin systems [33, 34, 35, 36, 37, 39, 40]. Beyond conventional MFTs, the corrections from quantum and thermal fluctuations are represented in terms of a series of $x^{n}$, where $x=H_{I}/(k_{B}T)$ and $H_{I}$ is part of the Hamiltonian describing the fluctuations [33, 34, 35]. Considering the series as a perturbative expansion ordered by the power $n$, $x$ can surpass the radius of convergence of the series at sufficiently low temperatures, which poses a challenge for convergence in LCE. This obstacle is better overcome by numerical LCE, where physical observables are calculated still in a basis as in LCE but not perturbatively in temperature [36, 37]. Then, it is the correlation length, instead of the temperature, that determines the convergence of the series. Consequently, numerical LCE can converge at a lower temperature than perturbative LCE. However, if long-range order develops at low temperature, and the correlation length becomes larger than the cut-off scale of the cluster sizes, the convergence is no longer guaranteed [38]. Resummation algorithms, such as Wynn’s algorithm, Brezinski’s algorithm [80], and Euler’s transformation [81], have been developed to accelerate the convergence of numerical LCE [38].
In DMMFT, the cluster partition is fixed at the beginning, and there are no summations over various configurations of the clusters. Although the size of the clusters plays a similar role in DMMFT and numerical LCE, setting the cut-off scale for the correlated fluctuations, DMMFT does not suffer from convergence problems even if long-range order develops at low temperatures because there are no explicit summations needed. Moreover, for practical applications, numerical LCE needs to handle a large number of large clusters of various sizes (often solved by ED), which can be computationally demanding. In this regard, DMMFT can outperform LCE in computational efficiency particularly at low temperatures.
@@ 4.2 Systems at Finite Temperatures
where the square bracket with a subscript $T$ indicates the thermal average. As discussed in Sec. 2.3, when $D_{\mathscr{H}_{\mathcal{C}^{\prime}_{c}}}=1$, DMMFT reduces to the CMFTs at zero temperature. It is straightforward to check that, since $[\rho_{c}^{\text{MF}}]_{T}\sim\exp[-\tilde{H}_{\text{MF}}^{(c)}/(k_{B}T)]$ when $D_{\mathscr{H}_{\mathcal{C}^{\prime}_{c}}}=1$, the extension of DMMFT according Eq.(28) also aligns with the conventional MFTs at finite temperatures.
However, it is worth noting that the thermal averaging in Eq.(28) does not fully capture all thermal fluctuations. Specifically, at low temperatures, it is often the gapless long wave length modes (Goldstone modes) that dominate the thermal fluctuations. However, due to the mean-field approximation inherent in DMMFT, these thermal fluctuations are frozen by the assumption of translational symmetry. Therefore, similar to all other mean-field methods, DMMFT tends to underestimate the impact of thermal fluctuations at finite temperatures.
@@ 4.3 Systems with Disorders
In real systems, the effects of disorders can be more complicated. For instance, there can be cluster of vacancies due to lower formation free energy, and the probability measure of the vacancies no longer takes the simple product form. Moreover, disorders can manifest as inhomogeneity of the samples, especially when scale of the probe is small, and the effects of disorders are not self-averaged in the experiments. Despite of the complications of disorders arising in real systems, DMMFT can, in principle, be adapted to include the disorder effects accordingly.
@@ 5 Conclusion
In this study, we introduce a novel mean-field approach, DMMFT, specifically designed to capture quantum fluctuations beyond conventional MFTs. Combining the strengths of DMRG and DMFT, DMMFT stands out as an efficient and unbiased method applicable to fermions, bosons, and spins, even in the presence of frustrations. A notable feature of DMMFT is its capability to discern topological phases enabled by its utilization of the reduced DM, as demonstrated with the example of the AKLT model in Sec. 3.1. Furthermore, DMMFT can achieve precision comparable to DMRG, even with only a small cluster, as demonstrated with the example of the AFHTL in Sec. 3.2. Additionally, DMMFT seamlessly extends to systems at finite temperatures and with disorders. In conclusion, DMMFT provides an effective tool for exploring phases exhibiting unconventional quantum orders. It can be particularly beneficial for investigating frustrated spin systems in high spatial dimensions.
@@ Appendix A Iterative Algorithm for Mean-Field Equations
Appendix B Pseudocodes for Heisenberg Model
In this Appendix section, we describe the implementation of DMMFT for the AFHTL that has been discussed in Sec. 3.2.
Figure 4: Extended cluster with periodic boundary conditions. The hexagons represent the Wigner-Seitz cells of the three-lattice clusters compatible with the symmetries of the underlying triangular lattice. The central cluster $c$ is highlighted in orange. Three neighboring clusters in $\mathcal{C}^{\prime}_{c}=\{c^{\prime}_{1},c^{\prime}_{2},c^{\prime}_{3}\}$ have been duplicated under periodic translations $\mathbf{a}_{i}$. Bonds depicted with solid lines indicate intra-cluster couplings contained in $H_{c}$, while dashed lines indicate inter-cluster couplings contained in $H_{c,\mathcal{C}^{\prime}_{c}}$. For visual clarity, not all coupling bonds are shown.
We choose a cluster partition of the triangular lattice with each cluster comprising three lattice sites as shown in Fig. 4. The hexagons represent the Wigner-Seitz cells of the three-lattice clusters compatible with the symmetries of the underlying triangular lattice. The central cluster $c$ is highlighted in orange. We choose an extended cluster $\bar{c}$ comprising of four three-lattice clusters $\{c;c^{\prime}_{1},c^{\prime}_{2},c^{\prime}_{3}\}$ with periodic boundary conditions. Fig. 4 shows the extended cluster in a periodically extended scheme, i.e., clusters in $\mathcal{C}^{\prime}_{c}$ have been duplicated under periodic translations $\mathbf{a}_{i}$. Bonds depicted in solid lines indicate intra-cluster couplings contained in $H_{c}$, while dashed lines indicate inter-cluster couplings contained in $H_{c,\mathcal{C}^{\prime}_{c}}$. For visual clarity, Fig. 4 has abbreviated some coupling bonds.
Now, we proceed to describe the pseudocodes of DMMFT.
First, we define the local degrees of freedom. For the AFHTL, on each site, there is a quantum spin of $S=\frac{1}{2}$. We define $S_{i}^{\alpha},\alpha=x,y,z$ and the identity operator for each site as follows.
1’# Code Block 1: define the physical site: S=1/2 #’
2% length of the spin
3slen = 1.0/2.0;
4% dimension of local Hilbert space
5dimHi = 2;
6% define local operators
7idiop = eye(dimHi);
8sxop = [0.0, 1.0; 1.0, 0.0]*slen;
9syop = [0.0, -1.0j; 1.0j, 0.0]*slen;
10szop = [1.0, 0.0; 0.0, -1.0]*slen;
Next, we define the local operators for the clusters shown in Fig. 4. These local operators for cluster $c$ are constructed from the site operators by tensor product, which are then represented in matrix form using the kron() function.
1’# Code Block 2: define local operators for cluster c #’
2% number of sites in c
3nsite = 3;
4% dimension of the cluster Hilbert space
5dimHc = dimHi^nsite;
6
7idcop = eye(dimHc);
8opclist = {idcop};
9
10for isite in c:
11 % construct siaop as tensor products of idiop’s saop
12 % for example: s1xop = kron(sxop,kron(idiop,idiop))
13 opclist.append(siaop);
14end for % isite
With local operators, the Hamiltonian $H_{c}$ containing intra-cluster exchange interactions and Zeeman couplings is constructed according to their definitions as follows.
1’# Code Block 3: construct Hamiltonian of intracluster couplings #’
2Hamc = zeros(dimHc);
3% add exchange interactions
4for ibond in c:
5 i,j = vertex(ibond);
6 Hamc = Hamc + sum_a Ja*siaop*sjaop;
7end for % ibond
8% add Zeeman couplings
9for isite in c:
10 Hamc = Hamc - sum_a ha*siaop;
11end for % ibond
Then, we define the operators for the extended cluster $\bar{c}$.
1’# Code Block 4: define operators for the extended cluster cbar #’
2% number of clusters in cbar
3ncbar = 4;
4% dimension of the cluster Hilbert space
5dimHcbar = dimHc*dimHcp^(ncbar-1); % cf. Eq.(10)
6
7idcbarop = eye(dimHcbar);
8opcbarlist = {idcbarop};
9
10% construct spin operators
11for isite in cbar:
12 % construct siabarop as tensor products of idcop’s siaop
13 % for example: s1xbarop = kron(s1xop,kron(idcop,kron(idcop,idcop)))
14 opcbarlist.append(siabarop);
15end for % isite
16
17% construct Hamiltonian operator with intra- and inter-cluster terms
[…]
Finally, we solve the mean-field equations self-consistently. We use the iterative method solving the mean-field equations, following the steps described in Append. A.
1’# Code Block 5: Iterative solver #’
2’> step 0: initialization <’
3rhoc = reduced DM from classical spin configurations
4
5% configure iteration process
6% learning rate
7alearn = 0.6;
8% cut-off dimension, n_c
9dimHcp = 4;
10
11% start iterations
12while (not converge):
13 ’> step 1: construct Projc [Eq.(13)] <’
14 evec_rhoc, eval_rhoc = eig(rhoc);
15 Projc = zeros(dimHc,nc);
16 for istate in largest n_c evals of rhoc:
17 Projc.append(evec_rhoc(istate));
18 end for % istate
19
20 ’> step 2: construct ProjCp [Eq.(14)] <’
21 for ic in C’:
22 ProjCp = kron(ProjCp,Projc)
23 end for % ic
24 ’> step 3: construct Proj [Eq.(10)] <’
25 Proj = kron(idcop,ProjCp);
26
27 ’> step 4: construct effective Hamiltonian [Eq.(11)] <’
28 Ham_MFeff = Proj’*Hamcbar*Proj;
29
30 ’> step 5: find ground state of Ham_MFeff <’
31 evec_Ham, eval_Ham = eig(Ham_MFeff);
32 phicbar = state in evec_Ham with lowest energy;
33 rhocbar = phicbar*phicbar’;
34
35 ’> step 6: calculate new rhoc [Eq.(9)] <’
36 rhoc_new = partially tracing rhocbar over {c’};
37
38 ’> step 0: update rhoc [Eq.(29)] <’
39 rhoc = (1-alearn)*rhoc + alearn*rhoc_new;
40 check convergence;
41
42end while % iteration
The pseudocode provided above serves only to demonstrate how to implement DMMFT step-by-step. Considering the memory and time efficiencies, many optimizations can be employed. For example, one may use the Lanczos algorithm in Code Block 5, line 14 and line 31, since not the full spectra of $\rho_{c}$ and $\tilde{H}_{\text{MF}}^{(c)}$ are used.